a car is moving at a speed of 40 m/s on a circular track of radius 400m. The speed is increasing at a rate 3m/s2 . the acceleration of the car is :
1.4m/s2
2.7m/s2
3.5m/s2
  1. 3m/s2

Grade:11

3 Answers

Hello
9 years ago
There are 2 components of acceleration.
  1. Towards center: V2/R = 40 x 40 /400 = 4 m/s2
  2. Tangent( orthogonal to center) = 3 m/s2
  3. Resultant of 1 and 2 above = 5 m/s2
Aaryan
6 years ago
Tangential acceleration =3m/s Radial acceleration = v^2/r =40*40/400 =4m/s Acceleration=√3^2+4^2 =√25=5m/s^2
Jethudan Charan
6 years ago
The only major noticeable point is that here not one acc is mentioned, radial or tangentially, so we will find the resultant of the two, radial (3m/s^2) and tangential (4 m /s^2)